A hypothetical alloy has a face-centered cubic unit cell. Its density is 10,470 kg/m3, and its atomic weight is 106.8 g/mole.

a) What is the atomic radius of the alloy in cm?

b) What is the volume of the single atom of the alloy in cm3?


1
Expert's answer
2021-04-21T04:04:14-0400

Convert the density into "g\/cm3"

D="10470kg\/m3"

D therefore= "10.47g\/cm3"


"d=(z\u00d7M)\/(a3\u00d7Na)"

For a FCC "Z=4"


Therefore,

10.47g/cm3= "(4\u00d7106.8g\/mol)\u00f7(a3\u00d76.022\u00d710^23)\n\na3=6.776\u00d710^-22."


But volume"= l3"

L= "3\u221avol"

"L=3\u221a6.776\u00d710^-22\n\nL=8.79\u00d710^-8cm"

"=8.79\u00d710^-8\u00d7(10^10\/10^2cm)"

=8.79∆


But, "C^2=a^2+a^2\n,C^2=2a^2"


"C^2=2\u00d7(8.79)^2"

"C^=\u221a154.53"

"C=12.43\u2206"


And,

"r=c\/4"

r=12.43/4

Ans for part( a)

"r=3.12" ans.



(B)

Vol of a single atom "V=4\/3\u03c0r^3"

"V=4\/3\u00d722\/7\u00d73.12^3"

"V=127.27cm^3" Ans.